MICROWAVE SPECTROSCOPY OF THE HCCS AND DCCS RADICALS IN THE VIBRONIC EXCITED STATES

\underline{JIAN TANG} AND SHUJI SAITO, Dept. of Molecular Structure, The Graduate University for Advanced Studies and Institute for Molecular Science, Okazaki 444, Japan.

The microwave spectra of the HCCS and DCCS radicals (X^{2}Π_{i}) in both CCS(ν_{5}) and H(D)CC(ν_{4}) bending vibrational states were observed in the frequency range of 160–400 GHz with a dc discharge of the C_{2}H_{2} (or C_{2}D_{2}) and CS_{2} gas mixture. The rotational transition series for the vibronic^{2}Δ_{5/2,3/2}(ν_{5}), μ, κ^{2}Σ(ν_{5}),^{2}Π_{3/2}(2ν_{5}), 2Δ_{5/2}(ν_{4}) and μ^{2}Σ(ν_{4}) states were assigned. The analysis of the effective molecular constants showed that the values of the γ_{eff} constants of the vibronic μ, κ^{2}Σ states were different from each other by 2–3 GHz, which could not be explained by the up-to-date Renner-Teller theory for a single bending mode. This abnormal Renner-Teller effect may be understood by a new vibronic interaction theory, which was developed by us to consider the cross vibronic interaction between the two bending modes.